# TriDi: Trilateral Diffusion of 3D Humans, Objects, and Interactions

### v2

@@ 1 Introduction
We propose TriDi, a unified 3D human-object interaction model capturing the joint distribution of humans, objects, and interactions. TriDi produces samples from every conditional distribution arising from the combination of three in addition to the joint distribution, giving rise to $72^{3}-1=$ 7 possible modes of operation, see Fig. 1.
@@ 2 Related Work
Modeling 3D HOI from the objects has been studied from diverse perspectives. At a macro scale, studying humans in the context of 3D scenes is prominent [26, 28, 100, 97, 31, 54, 74, 75, 91]. These works are instrumental for downstream tasks like synthetic dataset generation [4, 58, 82, 35]. Dynamic motions in scenes can be conditioned by object’s 3D location [8, 27], control points [98], milestones [60, 43], physical properties [99], or text descriptions [44, 66]. Such a high-level perspective on HOI should be complemented by modeling interactions on an object level. Hence,  [79, 68] synthesize the motion towards a static object, and [7, 45] focus on manipulation interactions. These works consider temporal sequences, which are demanding to capture, thus limiting the scaling beyond the settings seen at training time. Synthesizing hand-object interactions presents several challenges [17, 18], which originated specialized methods [12, 47, 88, 33, 93, 76, 49]. Producing accurate prediction raised the demand for hand-object refinement [53, 103, 69, 102], but those methods are limited by smaller objects and hand-held interactions. TriDi works with single frames, models contact beyond the hands, and thus supports human synthesis involving diverse objects.
@@
Contact is the physical medium of many human-object interactions. In practice, contact maps are a good proxy to promote realism [28, 56, 79]; however, their capture is often complicated by manual annotation [71] or the need for specialized hardware [5]. Contact is represented in a range of ways, e.g., as distances [15], proximity [97], or maps on the body [71] and the object [16]. Contact is often modeled on the hands [5, 6, 21], with recent works considering the full body [28, 71]. An alternative is to represent the interaction through text [15, 91, 66]. While more interpretable and controllable, this representation limits the possibility of spatial reasoning for the methods. In our work, we combine text and contact maps in a shared latent space, inheriting the advantage of both.
@@ 3 Background
Figure 2: TriDi Overview. TriDi is a Trilateral Diffusion for Human $\mathcal{H}$ (pose $\mathbf{\theta}_{\mathcal{H}}$, identity $\mathbf{\beta}_{\mathcal{H}}$, and 6-DoF global pose $\mathbf{g}_{\mathcal{H}}$), Object $\mathcal{O}$ (6-DoF global pose $\mathbf{g}_{\mathcal{O}}$) and Interaction $\mathcal{I}$ (Contact-Text latent $\mathbf{z}_{\mathcal{I}}$). In this figure the model is configured to sample $p({\color[rgb]{0.0625,0.4765625,0.07421875}{\mathcal{H}}},{\color[rgb]{0.0625,0.4765625,0.07421875}{\mathcal{I}}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}{\mathcal{O}}})$. One of the seven operating modes is chosen by adjusting the timestamp to be 00 for a given condition ($t^{\mathcal{O}}$ above) and $T$ for the desired prediction ($t^{\mathcal{H}}$ and $t^{\mathcal{I}}$ above), and supplying an object class condition (‘‘Table’’ above).
@@ 4.2 TriDi: Trilateral Diffusion for HOI
(^$\mathcal{F}(\hat{\mathcal{H}},\hat{\mathcal{O}},\hat{\mathcal{I}})=\sum\limits_{j\in|\mathbf{V}_{\mathcal{H}}|}|{\hat{\phi}_{\mathcal{I}}}{{}_{j}}\hat{\mathbf{d}}_{j}|,\vspace{-6pt}$ (11)
@@
Table 1: Quality of Generated Distribution. TriDi is the only one operating in all the modalities and shows better capability in covering data distribution, improving up to $47$%.
@@ 4.3 Training
To train our method, we must collect the contact maps $\phi_{\mathcal{I}}$ and the text descriptions $T_{\mathcal{I}}$, which are unavailable for many 3D HOI datasets. We define a simple automatic annotation procedure: for every training sample, we obtain the human-object distances $\mathbf{d}$ as described in Equation 12 and threshold them following [3] to obtain a binary contact map $\phi_{\mathcal{I}}$. We detect which of the $24$ body parts of the human template contains at least one vertex in contact, and we use this information to compose the labels following one of the predefined templates, e.g., "[parts] are in contact with [object]".
@@ 5.1 Comparison to one-way methods
We evaluate the quality of the generated distributions using three metrics and comparing the generated samples $g\in S_{g}$ with the reference ones $r\in S_{r}$ coming from the GT (with $|S_{g}|=|S_{r}|$), reporting the statistics across three sampling runs. The Coverage (COV[1] that measures the percentage of samples in $S_{r}$ that are matched with at least one sample from $S_{g}$ ($100$ indicates perfect overlap). The Minimum Matching Distance (MMD[1] measures the average distance of samples in $S_{r}$ to their closest neighbors in $S_{g}$, quantifying the misalignment of the distributions. The 1-nearest neighbor accuracy (1-NNA[84] measures the leave-one-out accuracy over the union of $S_{r}\cup S_{g}$; the optimal value is $50$, and it comprehensively assesses the quality of the distribution. To evaluate the Geometrical Consistency of Generation of humans, we report the Mean Per Joint Position Error (MPJPE) that measures the error in body joints in cm. We also report its value after applying Procustes Analysis (MPJPE-PA) [20], alleviating the effect of rotation and scale. For the object we employ the vertex-to-vertex ($\mathbf{E_{v2v}}$) and the object center ($\mathbf{E_{c}}$) errors, together with contact accuracy ($\mathbf{Acc_{cont}}$). For TriDi, we measure the error for the contact predicted directly by the method and the one calculated from the generated 3D HOI. We refer to Sup. Mat. for metrics’ rigorous definitions.
@@
TriDi is the first approach trained only once and modeling all the seven static HOI combinations. We compare with single-frame methods specialized in different modalities, posing a challenge to our general setup. We present a comparison with baselines in two modes, ${\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{H}},{\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{I}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{O}}$ and ${\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{O}},{\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{I}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{H}}$, as those are the most common ones and the only two that have existing methods working with static HOI. We present the further evaluation of other operating modes in Sup. Mat. For ${\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{H}},{\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{I}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{O}}$ we rely on GNet [68], while for ${\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{O}},{\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{I}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{H}}$ we choose state-of-the-art Object Pop-up [59] (ObjPOP). Since this latter is a regressive method, we provide a further baseline by substituting its object’s center MLP with a cVAE (ObjPOP [59]+cVAE). As in [59], we integrate a Nearest-Neighbor baseline (NN), which uses the condition modality to retrieve the closest example from the training set. For objects, the similarity is in terms of $\mathbf{g}_{\mathcal{O}}$, and for humans, it is the distance of body joints after centering the root. Additionally, to evaluate the benefits of the joint model, we train three variants of TriDi that work only in a single mode: s-TriDi-OI for ${\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{O}},{\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{I}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{H}}$, s-TriDi-HI for ${\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{H}},{\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{I}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{O}}$, and s-TriDi-HO for ${\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{H}},{\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{O}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{I}}$ (results in Sup. Mat.).
@@
For every method, we generate three samples for every example from the test sets of BEHAVE [3] and GRAB [67], and we compare the generated and the GT distributions. We consider cases when the condition modality is the object ${\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{H}},{\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{I}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{O}}$ or the human ${\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{O}},{\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{I}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{H}}$. We exclude NN and ObjPOP from this comparison since they do not have variance in prediction. In Tab.1, we report the metrics’ mean and variance. Despite being more general, TriDi outperforms the specialized baselines on all the metrics, improving up to 47%. The consistently higher COV and lower MMD indicate better diversity of the generated samples, while a 1-NNA close to $50$ suggests the generated samples are non-trivial. We report a qualitative comparison in Fig.4. Notably, TriDi performs better or on par with s-TriDi-HI and s-TriDi-OI , indicating that joint training benefits the generalization of the model. To compare the methods further, we conducted a user study that collected 40 responses. In summary, our method’s output is frequently preferred w.r.t. the baselines’ ones ($\sim$ 89% of the cases) and on par with the GT samples($\sim$ 52% of the cases); further details are in Sup. Mat.
@@
TriDi (Ours) $\mathbf{20.8}$ $\mathbf{12.3}$ $\nicefrac{{\underline{95.5}}}{{\mathbf{96.5}}}$ $28.0$ $\mathbf{15.3}$ $\nicefrac{{\mathbf{95.9}}}{{\underline{96.1}}}$
@@ 5.3 Applications
Populating scenes is interesting for several downstream tasks like those from AR/VR, or for synthetic data generations. Here, we demonstrate one possible way to populate scenes with TriDi. We first place a virtual object or a human on the ground of a scene, the publicly available HPS dataset[23], and then run TriDi to generate the complementary modality. Also, we can directly sample $p(\mathcal{H},\mathcal{O},\mathcal{I})$ and populate with both humans and objects. Results are shown in Fig. 6.
@@
Our method can also be used to reconstruct interactions from images indirectly. In Fig.7, we provide an example from the DAMON dataset of DECO [71]. DAMON provides contact annotations for images from the HOT [9] dataset along with SMPL parameters estimated by CLIFF [46]. Remarkably, TriDi generalizes to such cases despite not being trained on the DECO dataset. More examples are included in Sup. Mat.
@@
Perhaps the most widespread mechanism to generate sequences in animation is via keyframing. Here, we can use TriDi to automatically generate the keyframes of interaction and use an off-the-shelf in-betweener to generate a sequence. Given a 4D human sequence at 30 fps, we sample at 1 fps, generating the object and interpolate frames in between using slerp for angles and linear for translations. We provide an example in our Sup. Mat. video.
@@ 6 Conclusions
There are several exciting avenues for future work. We consider scaling beyond single human and object interaction a promising direction to enable the modeling of realistic social situations with increasing complexity. The recent advancements in data capturing [35, 95, 40] open up a possibility of blending scene and object conditioning, laying the foundation for advanced HOI models. As a data-driven method, TriDi is sensitive to the skewness of the data distribution, expressing more variety towards frequent objects. Although TriDi shows generalization to unseen geometries (e.g., chairs and stools), we do not expect it to support objects with significantly novel functionality (e.g., wheelchairs, bicycles, bowling balls).
@@ 8 Implementation details
The denoising network has a total of $15$M parameters, and it is trained end-to-end. We use a batch size of $1024$, a learning rate of $41e-$ 4 with a cosine scheduler, and warm up the training during the first $50$k steps. The parameters are optimized with AdamW [52]. We train for a total of $300$k steps. All the experiments are performed on a machine with RTX4090 GPU. The training of the model takes approximately $20$ hours. The contact encoder-decoder network with $1.7$M parameters is trained separately for $70$ epochs, converging on the same machine in $\sim 1$ hour. The inference for one example with diffusion guidance takes around $3.07$ seconds. Since TriDi works per-frame the inference can be majorly sped up using batching, e.g. inference time for $1024$ examples in one batch is $38.79$ s. All models are implemented in PyTorch [57] framework. Following [104] we convert all rotations ($\mathbf{\theta}_{\mathcal{H}},\mathbf{g}_{\mathcal{H}},\mathbf{g}_{\mathcal{O}}$) to 6-d representations before passing them to the network. We rely on blendify [24] for visualization.
We implement diffusion reconstruction guidance within the DDPM pipeline and apply it for the last $200$ out of $1000$ iterations of the denoising process with weight $\lambda=2.0$.

### v3

@@ 1 Introduction
We propose TriDi, a unified 3D human-object interaction model capturing the joint distribution of humans, objects, and interactions. TriDi produces samples from every conditional distribution arising from the combination of three in addition to the joint distribution, giving rise to $2^{3}-1=7$ possible modes of operation, see Fig. 1.
@@ 2 Related Work
Modeling 3D HOI from the objects has been studied from diverse perspectives. At a macro scale, studying humans in the context of 3D scenes is prominent [30, 32, 107, 104, 35, 59, 81, 82, 98]. These works are instrumental for downstream tasks like synthetic dataset generation [5, 63, 89, 39]. Dynamic motions in scenes can be conditioned by object’s 3D location [9, 31], control points [105], milestones [65, 47], physical properties [106, 69, 70], or text descriptions [48, 73]. Such a high-level perspective on HOI should be complemented by modeling interactions on an object level. Hence,  [86, 75] synthesize the motion towards a static object, and [8, 49] focus on manipulation interactions. These works consider temporal sequences, which are demanding to capture, thus limiting the scaling beyond the settings seen at training time. Synthesizing hand-object interactions presents several challenges [21, 22], which originated specialized methods [13, 51, 95, 37, 100, 83, 53]. Producing accurate prediction raised the demand for hand-object refinement [57, 111, 76, 110], but those methods are limited by smaller objects and hand-held interactions. TriDi works with single frames, models contact beyond the hands, and thus supports human synthesis involving diverse objects.
@@
Contact is the physical medium of many human-object interactions. In practice, contact maps are a good proxy to promote realism [32, 61, 86]; however, their capture is often complicated by manual annotation [78, 15] or the need for specialized hardware [6]. Contact is represented in a range of ways, e.g., as distances [18], proximity [104], or maps on the body [78] and the object [20, 19]. Contact is often modeled on the hands [6, 7, 25], with recent works considering the full body [32, 78, 19]. An alternative is to represent the interaction through text [18, 98, 73]. While more interpretable and controllable, this representation limits the possibility of spatial reasoning for the methods. In our work, we combine text and contact maps in a shared latent space, inheriting the advantage of both.
@@ 3 Background
Figure 2: TriDi Overview. TriDi is a Trilateral Diffusion for Human $\mathcal{H}$ (pose $\mathbf{\theta}_{\mathcal{H}}$, identity $\mathbf{\beta}_{\mathcal{H}}$, and 6-DoF global pose $\mathbf{g}_{\mathcal{H}}$), Object $\mathcal{O}$ (6-DoF global pose $\mathbf{g}_{\mathcal{O}}$) and Interaction $\mathcal{I}$ (Contact-Text latent $\mathbf{z}_{\mathcal{I}}$). In this figure the model is configured to sample $p({\color[rgb]{0.0625,0.4765625,0.07421875}{\mathcal{H}}},{\color[rgb]{0.0625,0.4765625,0.07421875}{\mathcal{I}}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}{\mathcal{O}}})$. One of the seven operating modes is chosen by adjusting the timestamp to be 0 for a given condition ($t^{\mathcal{O}}$ above) and $T$ for the desired prediction ($t^{\mathcal{H}}$ and $t^{\mathcal{I}}$ above), and supplying an object class condition (‘‘Table’’ above).
@@ 4.2 TriDi: Trilateral Diffusion for HOI
(^$\mathcal{F}(\hat{\mathcal{H}},\hat{\mathcal{O}},\hat{\mathcal{I}})=\sum\limits_{j\in|\mathbf{V}_{\mathcal{H}}|}|{\hat{\phi}_{\mathcal{I}}}{{}_{j}}\hat{\mathbf{d}}_{j}|,\vskip-6.0pt$ (11)
@@
Table 1: Quality of Generated Distribution. TriDi is the only one operating in all the modalities and shows better capability in covering data distribution, improving up to 474747%.
@@ 4.3 Training
To train our method, we must collect the contact maps $\phi_{\mathcal{I}}$ and the text descriptions $T_{\mathcal{I}}$, which are unavailable for many 3D HOI datasets. We define a simple automatic annotation procedure: for every training sample, we obtain the human-object distances $\mathbf{d}$ as described in Equation 12 and threshold them following [4] to obtain a binary contact map $\phi_{\mathcal{I}}$. We detect which of the 242424 body parts of the human template contains at least one vertex in contact, and we use this information to compose the labels following one of the predefined templates, e.g., "[parts] are in contact with [object]".
@@ 5.1 Comparison to one-way methods
We evaluate the quality of the generated distributions using three metrics and comparing the generated samples $g\in S_{g}$ with the reference ones $r\in S_{r}$ coming from the GT (with $|S_{g}|=|S_{r}|$), reporting the statistics across three sampling runs. The Coverage (COV[1] that measures the percentage of samples in $S_{r}$ that are matched with at least one sample from $S_{g}$ (100100100 indicates perfect overlap). The Minimum Matching Distance (MMD[1] measures the average distance of samples in $S_{r}$ to their closest neighbors in $S_{g}$, quantifying the misalignment of the distributions. The 1-nearest neighbor accuracy (1-NNA[91] measures the leave-one-out accuracy over the union of $S_{r}\cup S_{g}$; the optimal value is 505050, and it comprehensively assesses the quality of the distribution. To evaluate the Geometrical Consistency of Generation of humans, we report the Mean Per Joint Position Error (MPJPE) that measures the error in body joints in cm. We also report its value after applying Procustes Analysis (MPJPE-PA) [24], alleviating the effect of rotation and scale. For the object we employ the vertex-to-vertex ($\mathbf{E_{v2v}}$) and the object center ($\mathbf{E_{c}}$) errors, together with contact accuracy ($\mathbf{Acc_{cont}}$). For TriDi, we measure the error for the contact predicted directly by the method and the one calculated from the generated 3D HOI. We refer to Sup. Mat. for metrics’ rigorous definitions.
@@
TriDi is the first approach trained only once and modeling all the seven static HOI combinations. We compare with single-frame methods specialized in different modalities, posing a challenge to our general setup. We present a comparison with baselines in two modes, ${\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{H}},{\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{I}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{O}}$ and ${\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{O}},{\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{I}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{H}}$, as those are the most common ones and the only two that have existing methods working with static HOI. We present the further evaluation of other operating modes in Sup. Mat. For ${\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{H}},{\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{I}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{O}}$ we rely on GNet [75], while for ${\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{O}},{\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{I}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{H}}$ we choose state-of-the-art Object Pop-up [64] (ObjPOP). Since this latter is a regressive method, we provide a further baseline by substituting its object’s center MLP with a cVAE (ObjPOP [64]+cVAE). As in [64], we integrate a Nearest-Neighbor baseline (NN), which uses the condition modality to retrieve the closest example from the training set. For objects, the similarity is in terms of $\mathbf{g}_{\mathcal{O}}$, and for humans, it is the distance of root centered joints. Additionally, to evaluate the benefits of the joint model, we train three variants of TriDi that work only in a single mode: s-TriDi-OI for ${\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{O}},{\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{I}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{H}}$, s-TriDi-HI for ${\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{H}},{\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{I}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{O}}$, and s-TriDi-HO for ${\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{H}},{\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{O}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{I}}$ (results in Sup. Mat.).
@@
For every method, we generate three samples for every example from the test sets of BEHAVE [4] and GRAB [74], and we compare the generated and the GT distributions. We consider cases when the condition modality is the object ${\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{H}},{\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{I}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{O}}$ or the human ${\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{O}},{\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{I}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{H}}$. We exclude NN and ObjPOP from this comparison since they do not have variance in prediction. In Tab.1, we report the metrics’ mean and variance. Despite being more general, TriDi outperforms the specialized baselines on all the metrics, improving up to 47%. The consistently higher COV and lower MMD indicate better diversity of the generated samples, while a 1-NNA close to 505050 suggests the generated samples are non-trivial. We report a qualitative comparison in Fig.4. Notably, TriDi performs better or on par with s-TriDi-HI and s-TriDi-OI , indicating that joint training benefits the generalization of the model. To compare the methods further, we conducted a user study that collected 40 responses. In summary, our method’s output is frequently preferred w.r.t. the baselines’ ones ($\sim$ 89% of the cases) and on par with the GT samples ($\sim$ 52% of the cases); further details are in Sup. Mat.
@@
TriDi (Ours) $\mathbf{20.8}$ $\mathbf{12.3}$ $\nicefrac{{\underline{95.5}}}{{\mathbf{96.5}}}$ 28.028.028.0 $\mathbf{15.3}$ $\nicefrac{{\mathbf{95.9}}}{{\underline{96.1}}}$
@@ 5.3 Applications
Populating scenes is interesting for several downstream tasks in AR/VR, or for synthetic data generation. Here, we demonstrate one possible way to populate scenes with TriDi. We first place a virtual object or a human on the ground of a scene from the HPS dataset[27], and then run TriDi to generate the complementary modality. Also, we can directly sample $p(\mathcal{H},\mathcal{O},\mathcal{I})$ and populate with both humans and objects. Results are shown in Fig. 6.
@@
Our method can also be used to reconstruct interactions from images indirectly. In Fig.7, we provide an example from the DAMON dataset of DECO [78]. DAMON provides contact annotations for images from HOT [10] along with humans estimated by CLIFF [50]. Remarkably, TriDi generalizes to such cases despite not being trained on DECO. More examples are included in Sup. Mat. The requirement of a known 3D mesh can be relaxed by using recent image-to-mesh reconstruction methods such as SDFit [2].
@@
Perhaps the most widespread mechanism to generate sequences in animation is via keyframing. Here, we can use TriDi to automatically generate the keyframes of interaction and use an interpolation to generate a sequence. Given a 4D human sequence at 30 fps, we sample at 2 fps, generating the object and interpolate frames in between using slerp for angles and linear for translations. We provide an example in our Sup. Mat. video.
@@ 6 Conclusions
There are several exciting avenues for future work. We consider scaling beyond single human and object interaction a promising direction to enable the modeling of realistic social situations with increasing complexity. The recent advancements in data capturing [39, 102, 44, 58] open up a possibility of blending scene and object conditioning, laying the foundation for advanced HOI models. As a data-driven method, TriDi is sensitive to the skewness of the data distribution, expressing more variety towards frequent objects. Although TriDi shows generalization to unseen geometries (e.g., chairs and stools), we do not expect it to support objects with significantly novel functionality (e.g., wheelchairs, bicycles, bowling balls).
@@ 8 Implementation details
The denoising network has a total of  [^15] M parameters, and it is trained end-to-end. We use a batch size of 102410241024, a learning rate of $1e-4$ with a cosine scheduler, and warm up the training during the first 505050k steps. The parameters are optimized with AdamW [56]. We train for a total of 300300300k steps. All the experiments are performed on a machine with RTX4090 GPU. The training of the model takes approximately 202020 hours. The contact encoder-decoder network with 1.71.71.7M parameters is trained separately for 707070 epochs, converging on the same machine in 1$\sim$ 1∼ 1 hour. The inference for one example with diffusion guidance takes around 3.073.073.07 seconds. Since TriDi works per-frame the inference can be majorly sped up using batching, e.g. inference time for 102410241024 examples in one batch is 38.7938.7938.79 s. All models are implemented in PyTorch [62] framework. Following [112] we convert all rotations ($\mathbf{\theta}_{\mathcal{H}},\mathbf{g}_{\mathcal{H}},\mathbf{g}_{\mathcal{O}}$) to 6-d representations before passing them to the network. We rely on blendify [28] for visualization.
We implement diffusion reconstruction guidance within the DDPM pipeline and apply it for the last 200200200 out of 100010001000 iterations of the denoising process with weight $\lambda=2.0$.